If a sample of water contains 9.5mg of mgcl2 per litre calculate the hardness in terms of caco3
1
Expert's answer
2020-12-09T06:34:43-0500

m(Mg)=Ar(Mg)*m(MgCl2)/Mr(MgCl2)=24*9.5/95=2.4 mg

Hrdness= 2.4 mg/L